# Rotatable Antenna Aided Mixed Near-Field and Far-Field Communications in the Upper Mid-Band: Interference Analysis and Joint Optimization

### v2

@@ Abstract
In this paper, we propose to leverage rotatable antennas (RAs) for improving the communication performance in mixed near-field and far-field communication systems by exploiting a new spatial degree-of-freedom (DoF) offered by antenna rotation to mitigate complex near-field interference and mixed-field interference. Specifically, we investigate a modular RA-enabled mixed-field downlink communication system, where a base station (BS) consisting of multiple RA subarrays communicates with multiple near-field users in the presence of several legacy far-field users. We formulate an optimization problem to maximize the sum-rate of the near-field users by jointly optimizing the power allocation and rotation angles of all subarrays at the BS. To gain useful insights into the effect of RAs on mixed-field communications, we first analyze a special case where all subarrays share the same rotation angle and obtain closed-form expressions for the rotation-aware normalized near-field interference and the rotation-aware normalized mixed-field interference using the Fresnel integrals. We then analytically reveal that array rotation effectively suppresses both interference types, thereby significantly enhancing mixed-field communication performance. For the general case involving subarray-wise rotation, we propose an efficient double-layer algorithm to obtain a high-quality solution, where the inner layer optimizes power allocation using the successive convex approximation (SCA) technique, while the outer layer determines the rotation angles of all subarrays via particle swarm optimization (PSO). Finally, numerical results highlight the significant performance gains achieved by RAs over conventional fixed-antenna systems and demonstrate the effectiveness of our developed joint design compared to benchmark schemes.
Index Terms:
Rotatable antenna (RA), upper mid-band, mixed near-field and far-field communications, interference analysis.
@@ I Introduction
To fulfill the ever-growing demands of future sixth-generation (6G) wireless networks for ultra-high data rates, hyper-reliability, and extremely low latency [1, 2, 3], communication systems must undergo substantial evolution. At the core of this transformation, and critical to shaping 6G and potentially next-generation (NextG) standards, are advancements in multiple-input multiple-output (MIMO) technologies, particularly the deployment of extremely large-scale MIMO (XL-MIMO), along with the strategic utilization of new spectrum bands such as frequency range $3$ (FR3), which spans $7$ GHz to $24$ GHz [4]. These key enabling technologies are anticipated to deliver substantial benefits for 6G, enhancing end-user experiences through improved throughput, extended coverage, and greater energy efficiency. Consequently, they are expected to significantly influence future standardization efforts within bodies like the 3rd Generation Partnership Project (3GPP) [5]. These combined thrusts of XL-MIMO and new spectrum utilization are essential for achieving unprecedented spectral efficiency and spatial resolution [6, 7]. More importantly, this combination fundamentally reshapes the electromagnetic properties of the wireless environment. The synergistic impact of MIMO’s evolution towards XL-MIMO, together with the adoption of new spectrum like FR3, signifies that the transition from conventional far-field communications (planar-wave propagation) to near-field communications (spherical-wave propagation) [8, 9], and ultimately to the more general mixed near-field and far-field communications [10], is no longer merely a theoretical outcome. Instead, it emerges as an indispensable and practical consideration for air interface design in 6G and NextG systems. For instance, in an XL-MIMO system where a base station (BS) is equipped with $128$ antennas of half-wavelength inter-antenna spacing and operates at $24$ GHz, the Rayleigh distance is approximately $88$ meter (m). Consequently, considering the typical cell sizes in fifth-generation (5G) networks, users are likely to be located in both the near-field and far-field regions relative to the BS. This increasingly prevalent mixed-field scenario introduces complex design challenges that demand careful consideration in standardization efforts, including mixed-field interference management [10] and mixed-field integrated sensing and communications (ISAC) [11]. Moreover, understanding and leveraging these mixed-field characteristics will be crucial for realizing novel applications and ensuring their robust operation through effective standardization, such as mixed-field user coexistence [12] as well as mixed-field simultaneous wireless information and power transfer [13]. Addressing these aspects is fundamental to unlocking the full potential of 6G, particularly in the evolution of MIMO systems within FR3 and potentially higher frequency bands.
@@
More specifically, unlike purely near-field or far-field communications, prior work [10] has highlighted a key characteristic inherent to mixed-field communications: a near-field user experiences significant interference from the beam intended for a far-field user, even when positioned at a different spatial angle. This interference stems from the energy-spread effect caused by the mismatch between the far-field beamforming vector and the near-field steering vector. Such a mismatch leads to strong correlations across a wide angular domain, thereby inevitably resulting in complex mixed-field interference issues. This interference, combined with pure near-field and far-field inter-user interference, significantly complicates the design for mixed-field communication systems. Moreover, existing work on mitigating such interference issues is still in its infancy and has not been thoroughly investigated.
@@
Movable antennas (MAs)/fluid antennas [14, 15, 16, 17, 18, 19, 20] have recently emerged as a promising solution for interference management by exploiting the degree-of-freedom (DoF) offered by flexible antenna movement. Specifically, the authors in [19] demonstrated that by properly designing the antenna position, full array gain can be achieved in the desired direction while completely nullifying interference in all undesired directions. Subsequently, this study was further extended in [20] to multiple desired directions, which suggested maximizing the beamforming gain across signal directions while constraining the maximum interference power in undesired directions. However, it is important to note that the aforementioned works are designed specifically for interference mitigation in far-field communications, which may not be applicable to mixed-field communication scenarios. In such scenarios, a complex interplay of near-field inter-user interference, far-field inter-user interference, and mixed-field inter-user interference coexists, necessitating advanced approaches.
@@
To fully exploit the six-dimensional (6D) spatial DoFs, the concept of the 6D movable antenna (MA) has recently been introduced, allowing flexible adjustment of both the three-dimensional (3D) position and 3D rotation of antennas/arrays [21, 22, 23, 24]. Although this additional flexibility significantly enhances adaptability to the wireless environment, it also incurs substantial design complexity and implementation cost due to the intricate mechanical constraints associated with antenna movements. To alleviate these challenges, rotatable antennas (RAs) have been proposed as a simplified and cost-effective implementation of 6DMA [25, 26, 27, 28, 29, 30]. RAs can be considered as a special case of 6DMA by maintaining fixed antenna positions while allowing antenna rotation. However, RAs greatly reduce hardware complexity yet still offer the capability to reshape radiation patterns and enable flexible coverage. Initial studies have explored the integration of RAs into various wireless systems. For instance, in [26], a new RA model was proposed to enhance the communication performance by independently adjusting each antenna’s boresight direction to achieve a desired array gain pattern. Later on, in [27], RAs were integrated into the THz wideband systems to mitigate the beam-squint effects, capitalizing on the absence of beam squint along the boresight direction. Furthermore, the application of RAs in emerging communication scenarios has attracted significant interest. For example, in [28], RAs are utilized to secure information transmission between a BS and a legitimate user against multiple eavesdroppers, while [29] and [30] leveraged RAs to enhance ISAC performance in both single-BS and cooperative BS scenarios. Despite these advances, the potential of RAs in mixed-field communication systems remains largely unexplored, particularly regarding their ability to mitigate the unique interference challenges inherent to mixed-field scenarios.
Motivated by the above, we consider in this paper a new RA-enabled mixed near-field and far-field communication system and investigate the performance gains achievable through RAs. Specifically, we consider a scenario where a modular RA-array, comprising multiple subarrays, is deployed at a BS to serve multiple near-field users while coexisting with multiple legacy far-field users supported by preconfigured spatial beams, as illustrated in Fig. 1. We focus on the joint design of power allocation for near-field users and rotation angles of all RA subarrays at the BS, where the BS is assumed to simultaneously steer multiple beams towards near-field users by leveraging the appealing near-field beam-focusing property, along with appropriately designed power allocation. The main contributions are summarized as follows.
  • First, to the best of our knowledge, we are the first to propose leveraging RAs for improving the communication performance in mixed near-field and far-field communication systems. We formulate an optimization problem to maximize the sum-rate of near-field users by jointly optimizing the power allocation for near-field users and the rotation angles of all RA subarrays, subject to the constraints on the BS transmit power and the admissible rotation range of each subarray.
  • Second, to shed important insights into the effect of RAs on mixed-field communication performance, we first analyze a special case where all subarrays share the same rotation angle. In particular, we derive closed-form expressions for the rotation-aware normalized near-field interference and the rotation-aware normalized mixed-field interference using the Fresnel integrals. We then analytically demonstrate that array rotation effectively suppresses both near-field and mixed-field interference, thereby significantly enhancing mixed-field communication performance.
  • Third, for the general case involving subarray-wise rotation, the resultant problem corresponds to non-convex optimization with highly coupled variables. To tackle this, we devise an efficient double-layer algorithm to obtain a high-quality solution. Specifically, the proposed approach consists of two layers: the inner layer optimizes the power allocation via the successive convex approximation (SCA), while the outer layer determines the rotation angles of all subarrays using the particle swarm optimization (PSO).
  • Finally, numerical results highlight the significant benefits of integrating RAs into mixed-field communication systems, as well as validate the effectiveness of the proposed joint design. In particular, it reveals that: 1) RA-enabled systems significantly outperform conventional fixed-antenna systems in mixed-field communication scenarios; 2) subarray-wise rotation offers superior performance compared to uniform array rotation due to greater design flexibility; and 3) the proposed joint design substantially surpasses benchmark schemes that only optimize either power allocation or rotation angles of RA subarrays.
@@
Notations: Bold lower-case letters denote the column vectors. For a vector $\mathbf{x}$, $\mathbf{x}]_{i}$ indicates its $i$-th element. $\mathbb{C}$ represents the set of complex numbers. The superscripts $(\cdot)^{T}$ and $(\cdot)^{H}$ represent the transpose and the conjugate transpose, respectively. $\mathcal{CN}(0,\sigma^{2})$ denotes the zero-mean circularly symmetric complex Gaussian (CSCG) distribution with variance $\sigma^{2}$. The notation $|\cdot|$ denotes the absolute value of a real number or the cardinality of a set. The matrix $\mathbf{I}_{N}$ represents the identity matrix of size $N$. $\mathcal{O}(\cdot)$ denotes the standard big-O notation.
@@ II System Model and Problem Formulation
As illustrated in Fig. 1, we consider a mixed near-field and far-field downlink communication system enabled by modular RAs. The system comprises a BS equipped with a uniform linear array (ULA) of $N=2\tilde{N}+1$ antennas, which serves $K$ single-antenna near-field users indexed by $\mathcal{K}=\{1,2,\ldots,K\}$ in the presence of $M$ legacy single-antenna far-field users indexed by $\mathcal{M}=\{1,2,\ldots,M\}$. The ULA has an aperture of D=(N1)dD=(N-1)d, where $d=\frac{\lambda}{2}$ represents the inter-element spacing, with $\lambda$ denoting the carrier wavelength. The distinction between near-field and far-field users is determined by the Rayleigh distance, defined as $Z_{\rm Ray}=\frac{2D^{2}}{\lambda}$. Specifically, near-field users are located at distances smaller than $Z_{\rm Ray}$ from the BS, while far-field users are positioned beyond this threshold.
@@ II-A1 Modular RA architecture
The ULA is divided into $Q$ subarrays, indexed by $\mathcal{Q}=\{-\tilde{Q},\ldots,\tilde{Q}\}$, with each subarray consisting of $\bar{N}$ antennas. We assume that $\bar{N}$ is an odd number and the antennas within each subarray $q\in\mathcal{Q}$ are indexed by $\bar{n}\in\mathcal{\bar{N}}=\{-\frac{\bar{N}-1}{2},\ldots,\frac{\bar{N}-1}{2}\}$. [^1] We assume that each subarray comprises $\bar{N}=\lfloor\frac{N}{Q}\rfloor$ antennas, with the remaining antennas allocated to the leftmost and rightmost subarrays, respectively. In particular, $Q$ subarrays are connected to a common smart controller [26], wherein each subarray $q\in\mathcal{Q}$ can be electrically rotated around the $x$-axis by an angle $\phi_{q}$ to facilitate the subarray-wise rotation for improving the mixed-field communication performance [29]. [^2] In this paper, to theoretically evaluate the performance gains of RAs, we assume that the BS employs a ULA with subarray-wise 1D rotation [29]. Moreover, the implementation of ULA rotation requires minimal mechanical components in practice, enhancing its practicality for real-world deployment. Notably, the extension to more general cases, including uniform planar array (UPA) and/or three-dimensional (3D) rotation, is straightforward, which will be discussed in detail in Section III-D. The position of the $\bar{n}$-th antenna, before rotation, is defined relative to the subarray’s center at ${c}_{q}=q\bar{N}$ along the $x$-axis. After rotation by $\phi_{q}$, the antenna’s coordinate becomes $\mathbf{w}_{q,\bar{n}}=[\delta_{q,\bar{n}}\cos{\phi_{q}},\delta_{q,\bar{n}}\sin{\phi_{q}}]^{T}$, where $\delta_{q,\bar{n}}={c}_{q}+\bar{n}d$ $\forall{q\in\mathcal{Q}}$ and $\bar{n}\in\mathcal{\bar{N}}$. Based on the above, the channel models for near-field and far-field users accounting for the subarray-wise rotation are described as follows.

### v4

@@ Abstract
In this paper, we propose to leverage rotatable antennas (RAs) for improving the communication performance in mixed near-field and far-field communication systems by exploiting a new spatial degree-of-freedom (DoF) offered by antenna rotation to mitigate complex near-field interference and mixed-field interference. Specifically, we investigate an RA-enabled mixed-field downlink communication system with a modular architecture, where the base station (BS) is composed of multiple RA subarrays serving multiple near-field users in the presence of legacy far-field users. We formulate an optimization problem to maximize the sum-rate of the near-field users by jointly optimizing the power allocation and rotation angles of all subarrays at the BS. To gain useful insights into the effect of RAs on mixed-field communications, we first analyze a special case where all subarrays share the same rotation angle and obtain closed-form expressions for the rotation-dependent normalized near-field interference and the rotation-dependent normalized mixed-field interference using the Fresnel integrals. We then analytically reveal that array rotation effectively suppresses both interference types, thereby significantly enhancing mixed-field communication performance. For the general case involving subarray-wise rotation, we propose an efficient double-layer algorithm to obtain a high-quality solution, where the inner layer optimizes power allocation using the successive convex approximation (SCA), while the outer layer determines the rotation angles of all subarrays via particle swarm optimization (PSO). Finally, numerical results highlight the significant performance gains achieved by RAs over conventional fixed-antenna systems and demonstrate the effectiveness of the developed joint design compared to benchmark schemes.
@@ I Introduction
To fulfill the ever-growing demands of future sixth-generation (6G) wireless networks for ultra-high data rates, hyper-reliability, and extremely low latency [1, 2], communication systems must undergo substantial evolution. At the core of this transformation, and critical to shaping 6G and potentially next-generation (NextG) standards, are advancements in multiple-input multiple-output (MIMO) technologies, particularly the deployment of extremely large-scale MIMO (XL-MIMO), along with the strategic utilization of new spectrum bands such as frequency range $3$ (FR3), which spans $7$ GHz to $24$ GHz [3]. These key enabling technologies are anticipated to deliver substantial benefits for 6G, enhancing end-user experiences through improved throughput, extended coverage, and higher energy efficiency. Consequently, they are expected to significantly influence future standardization efforts within bodies like the 3rd Generation Partnership Project (3GPP) [4]. These combined thrusts of XL-MIMO and new spectrum utilization are essential for achieving unprecedented spectral efficiency and spatial resolution [5, 6]. More importantly, this integration fundamentally reshapes the electromagnetic properties of the wireless environment. The synergistic impact of MIMO’s evolution towards XL-MIMO, together with the adoption of new spectrum like FR3, signifies that the transition from conventional far-field communications (planar-wave propagation) to near-field communications (spherical-wave propagation) [7, 8], and ultimately to the more general mixed near-field and far-field communications [9], is no longer merely a theoretical assumption. Instead, it emerges as an indispensable and practical consideration for air interface design in 6G and NextG systems. For instance, in an XL-MIMO system where a base station (BS) is equipped with $128$ antennas at half-wavelength inter-element spacing operating at $24$ GHz, the Rayleigh distance is approximately $100$ meter (m). Consequently, considering the typical cell sizes in fifth-generation (5G) networks, there are likely users in both the near-field and far-field regions of the BS. This increasingly prevalent mixed-field scenario introduces complex design challenges that demand careful consideration in standardization efforts, including mixed-field interference management [9] and mixed-field integrated sensing and communications (ISAC) [10]. Moreover, understanding and leveraging these mixed-field characteristics will be crucial for realizing novel applications and ensuring their robust operation, facilitating mixed-field user coexistence [11] as well as mixed-field simultaneous wireless information and power transfer [12].
@@
More specifically, unlike purely near-field or far-field communications, prior work [9] has highlighted a key characteristic inherent to mixed-field communications: a near-field user experiences significant interference from beams intended for far-field users, even when positioned at a different spatial angle. This interference, referred to as mixed near-field and far-field interference, arises from the energy-spread effect caused by the mismatch between the far-field beamforming vector and the near-field steering vector. This mismatch leads to strong correlations across a wide angular domain, thereby inevitably resulting in complex mixed-field interference challenges. This interference, combined with pure near-field and far-field inter-user interference, significantly complicates the design of mixed-field communication systems. Moreover, existing work on mitigating such interference is still in its infancy, and the related challenges have not been thoroughly investigated.
@@
Movable antennas (MAs)/fluid antennas [13, 14, 15, 16, 17, 18, 19, 20] have recently emerged as a promising solution for interference management by exploiting the degrees-of-freedom (DoFs) offered by flexible antenna movement. Specifically, the authors in [19] demonstrated that by properly designing the antenna position, full array gain can be achieved in the desired direction while completely nullifying interference in all undesired directions. Subsequently, this study was further extended in [20] to multiple desired directions, where the beamforming gain across signal directions is maximized while the maximum interference power in undesired directions is constrained. However, the aforementioned works specifically tackled interference mitigation for far-field communications and may not be applicable to mixed-field communication scenarios. In such scenarios, a complex interplay of near-field inter-user interference, far-field inter-user interference, and mixed-field inter-user interference exists, necessitating advanced interference mitigation approaches.
@@
To fully exploit the six-dimensional (6D) spatial DoFs, the concept of 6D movable antennas (6DMAs) has recently been introduced, allowing flexible adjustment of both the three-dimensional (3D) position and 3D rotation of antennas/arrays [21, 22, 23, 24]. Although this additional flexibility significantly enhances adaptability to the wireless environment, it also incurs substantial design complexity and implementation cost due to the intricate mechanical constraints associated with antenna movement. To alleviate these challenges, rotatable antennas (RAs) have been proposed as a simplified and cost-effective 6DMA implementation [25, 26, 27, 28, 29, 30]. RAs can be considered as a special case of 6DMAs, maintaining fixed antenna positions while allowing antenna rotation. Accordingly, RAs greatly reduce hardware complexity yet still offer the capability to reshape radiation patterns and enable flexible coverage. Initial studies have explored the integration of RAs into various wireless systems. For instance, in [26], a new RA model was proposed to enhance the communication performance by independently adjusting each antenna’s boresight direction to achieve a desired array gain pattern. In addition, [28] employs RAs to enable secure data transmission between a BS and a legitimate user in the presence of multiple eavesdroppers, while the authors in [29] and [30] optimize the orientation of RA array to enhance ISAC performance in single-BS and cooperative BS scenarios, respectively. Despite these advances, the potential of RAs in mixed-field communication systems remains largely unexplored, and their ability to tackle the unique interference challenges inherent to mixed-field scenarios has not been thoroughly studied yet. Moreover, the existing literature on RAs mainly focuses on optimizing rotation angles either element-wise or array-wise, which fails to strike a balance between design flexibility and practical implementability. This limitation calls for a more flexible architecture that can better navigate this trade-off.
@@
Motivated by the above, in this paper, we consider a new modular RA-enabled mixed near-field and far-field communication system and investigate the performance gains achievable through RAs. Specifically, we consider a scenario where a modular RA-array, comprising multiple subarrays, is deployed at a BS to serve multiple near-field users in addition to multiple legacy far-field users supported by preconfigured spatial beams, as illustrated in Fig. 1. We focus on the joint design of the power allocation for the near-field users and the rotation angles of all RA subarrays at the BS, where the BS is assumed to simultaneously steer multiple beams towards the near-field users by leveraging the appealing near-field beam-focusing property. The main contributions of the paper are summarized as follows.
  • First, to the best of our knowledge, we are the first to propose leveraging RAs for improving the communication performance in mixed near-field and far-field communication systems. We formulate an optimization problem to maximize the sum-rate of the near-field users by jointly optimizing the power allocation for the near-field users and the rotation angles of all RA subarrays, subject to constraints on the BS transmit power and the admissible rotation range of each subarray.
  • Second, to shed light onto the impact of RAs on mixed-field communication performance, we first analyze a special case where all subarrays share the same rotation angle. For this case, we derive closed-form expressions for the rotation-dependent normalized near-field interference and the rotation-dependent normalized mixed-field interference using the Fresnel integrals. We then analytically demonstrate that array rotation effectively suppresses both near-field and mixed-field interference, thereby significantly enhancing mixed-field communication performance.
  • Third, for the general case of subarray-wise rotation, a non-convex optimization problem with highly coupled variables is obtained. To tackle this, we devise an efficient double-layer algorithm to obtain a high-quality solution. Specifically, the proposed approach consists of two layers: the inner layer optimizes the power allocation via successive convex approximation (SCA), while the outer layer determines the rotation angles of the subarrays using particle swarm optimization (PSO).
  • Finally, our numerical results highlight the significant benefits of integrating RAs into mixed-field communication systems and validate the effectiveness of the proposed joint design. In particular, we reveal that: 1) RA-enabled systems significantly outperform conventional fixed-antenna systems in mixed-field communication scenarios; 2) subarray-wise rotation offers superior performance compared to uniform array rotation due to greater design flexibility; and 3) the proposed joint design substantially surpasses benchmark schemes that only optimize either power allocation or rotation angles of RA subarrays.
Notations: Bold lower-case letters denote column vectors. For a vector $\mathbf{x}$, $\mathbf{x}]_{i}$ indicates its $i$-th element. $\mathbb{C}$ represents the set of complex numbers. $\lfloor x\rfloor$ denotes the floor of $x$, i.e., the largest integer less than or equal to $x$. The superscripts $(\cdot)^{T}$ and $(\cdot)^{H}$ represent the transpose and the conjugate transpose, respectively. $\mathcal{CN}(0,\sigma^{2})$ denotes the zero-mean circularly symmetric complex Gaussian (CSCG) distribution with variance $\sigma^{2}$. The notation $|\cdot|$ denotes the absolute value of a real number or the cardinality of a set. Matrix $\mathbf{I}_{N}$ represents the identity matrix of size $N$. $\mathcal{O}(\cdot)$ denotes the standard big-O notation.
@@ II System Model and Problem Formulation
As illustrated in Fig. 1, we consider a mixed-field downlink communication system enabled by modular RAs. The system comprises a BS equipped with a uniform linear array (ULA) of $N=2\tilde{N}+1$ antennas, which serves $K$ single-antenna near-field users indexed by $\mathcal{K}=\{1,2,\ldots,K\}$ in the presence of $M$ legacy single-antenna far-field users indexed by $\mathcal{M}=\{1,2,\ldots,M\}$. The ULA has an aperture of D=(N1)dD=(N-1)d, where $d=\frac{\lambda}{2}$ represents the inter-element spacing, with $\lambda$ denoting the carrier wavelength. The distinction between near-field and far-field users is determined by the Rayleigh distance, defined as $Z_{\rm Ray}=\frac{2D^{2}}{\lambda}$. Specifically, near-field users are located at distances smaller than $Z_{\rm Ray}$ from the BS, while far-field users are positioned beyond this threshold.
@@ II-A1 Modular RA architecture
The ULA is divided into $Q$ subarrays, indexed by $\mathcal{Q}=\{-\tilde{Q},\ldots,\tilde{Q}\}$, with each subarray consisting of $\bar{N}$ antennas. We assume that $\bar{N}$ is an odd number and the antennas within each subarray $q\in\mathcal{Q}$ are indexed by $\bar{n}\in\mathcal{\bar{N}}=\{-\frac{\bar{N}-1}{2},\ldots,\frac{\bar{N}-1}{2}\}$. [^1] We assume that each subarray comprises $\bar{N}=\lfloor\frac{N}{Q}\rfloor$ antennas, with the remaining antennas allocated to the leftmost and rightmost subarrays, respectively. In particular, $Q$ subarrays are connected to a common smart controller [26], wherein each subarray $q\in\mathcal{Q}$ can be electrically rotated around the $x$-axis by an angle $\phi_{q}$ to facilitate the subarray-wise rotation for improving the mixed-field communication performance [29]. [^2] In this paper, to theoretically evaluate the performance gains of RAs, we assume that the BS employs a ULA with subarray-wise 1D rotation [29]. Moreover, the implementation of ULA rotation requires minimal mechanical components in practice, enhancing its practicality for real-world deployment. Notably, the extension to more general cases, including uniform planar array (UPA) and/or three-dimensional (3D) rotation, is straightforward, which will be discussed in detail in Section III-D. The position of the $\bar{n}$-th antenna, before rotation, is defined relative to the subarray’s center at ${c}_{q}=q\bar{N}d$ along the $x$-axis. After rotation by $\phi_{q}$, the antenna’s coordinate become $\mathbf{w}_{q,\bar{n}}=[c_{q}+\bar{n}d\cos{\phi_{q}},\bar{n}d\sin{\phi_{q}}]^{T}$, $\forall{q\in\mathcal{Q}}$ and $\bar{n}\in\mathcal{\bar{N}}$. Based on the above, the channel models for near-field and far-field users accounting for the subarray-wise rotation are described as follows.